vineri, 21 ianuarie 2011

Commander Check-in Report Day 06 RoMars2011

Log Book for January 20, 2011
Commander's Check-In
Virgiliu Pop Reporting

Nikon pe Marte
Crew photos for January 20, 2011.

Crew Physical Status: Excellent

Time Departed/Returned from EVA: 13:00-17:30

Brief Narrative of Field Mission Results: Soil samples collected.

EVA Data/Interpretations: EVA number 3; Samples to be analyzed.

Engineering/Hab Maintenance: We started to remove the old microphones from the spacesuit helmets; ATV tires got inflated.

Report Transmission Schedule: Intended by 8 PM.

Plans for Tomorrow: Will be decided in the morning.

Inventory: One extra mouse.

Support Requested: Pipes need to be fixed.

Miscellaneous: Third SIM day.


Trimiteți un comentariu